# RepairLLaMA: Efficient Representations and Fine-Tuned Adapters for Program Repair

### v3

@@
Abstract.
Automated Program Repair (APR) has evolved significantly with the advent of Large Language Models (LLMs). Fine-tuning LLMs for program repair is a recent avenue of research, with many dimensions which have not been explored. Existing work mostly fine-tunes LLMs with naive code representations and is fundamentally limited in its ability to fine-tune larger LLMs. To address this problem, we propose RepairLLaMA, a novel program repair approach that combines 1) code representations for APR and 2) the state-of-the-art parameter-efficient LLM fine-tuning technique called LoRA. This results in RepairLLaMA producing a highly effective ‘program repair adapter’ for fixing bugs with language models. Our experiments demonstrate the validity of both concepts. First, fine-tuning adapters with program repair specific code representations enables the model to use meaningful repair signals. Second, parameter-efficient fine-tuning helps fine-tuning to converge and contributes to the effectiveness of the repair adapter to fix data-points outside the fine-tuning data distribution. Overall, RepairLLaMA correctly fixes 125 Defects4J v2 and 82 HumanEval-Java bugs, outperforming all baselines.
@@ Abstract.
**footnotetext: These authors contributed equally to this work.
1. Introduction
@@
Automated program repair (APR) (Monperrus, 2018) aims at automatically fixing a software bug without human intervention. Learning-based repair (Chen et al., 2019; Zhu et al., 2021; Jiang et al., 2021a; Ye et al., 2022; Xia and Zhang, 2022; Jiang et al., 2023; Xia et al., 2023; Wang et al., 2023b) has become the mainstream solution to this problem due to the powerful ability of deep neural networks to learn complex bug fix patterns. Clearly, large language models (LLMs), pre-trained on vast amounts of data, have pushed learning-based repair to the next frontier (Jiang et al., 2023; Xia et al., 2022). In program repair, LLMs have been mostly used with prompt engineering (Joshi et al., 2023; Xia and Zhang, 2023), and recently, a line of work around fine-tuning has emerged (Jiang et al., 2021a; Yuan et al., 2022; Jiang et al., 2023; Wang et al., 2023b).
@@
Fine-tuning LLMs for program repair is complex. Early work simply refines the network weights based on additional fine-tuning data. However, this kind of fine-tuning is rather primitive and suffers from two significant drawbacks. First, fine-tuning is also known to be able to adapt the input/output representations of the data under study (Chen et al., 2021a). In the context of program repair, there is an opportunity to fine-tune with code representations that maximize downstream task performance, that is, repair performance. In particular, previous work overlooks the realistic representation of fault localization in the input. Second, previous work considered the most basic fine-tuning technique, which is full-parameter fine-tuning. As LLMs increase in size (Lu et al., 2023), full-parameter fine-tuning poses important overfitting problems when fine-tuning data is limited, which is typically the case in program repair. In this paper, we address the problem of devising efficient fine-tuning techniques (Hu et al., 2021) for program repair, with a focus on code representations and adapters.
@@
We propose RepairLLaMA, a new program repair approach that combines realistic repair-specific code representations with parameter-efficient fine-tuning. First, RepairLLaMA’s code representations incorporate fault localization signals and are designed to support multi-location bugs. Second, RepairLLaMA utilizes Low-Rank Adaption (LoRA), a state-of-the-art parameter-efficient fine-tuning technique, to train a much smaller repair adapter (when compared to the full LLM) that adapts the LLM for program repair while helping prevent overfitting. As we will demonstrate in this paper, the concept of repair adapter is novel and potent.
@@
Our experimental results validate RepairLLaMA’s core designs. First, RepairLLaMA achieves state-of-the-art performance in two benchmarks, correctly fixing 125 Defects4J v2 (Just et al., 2014) bugs and 82 bugs on recently proposed HumanEval-Java (Jiang et al., 2023), which boosts internal and external validity. The experiments show that the devised code representations with repair signals allow the LLM to synthesize patches more effectively than with a naive code-only representation. Also, RepairLLaMA clearly outperforms non-fine-tuned baselines, incl. GPT-4. Moreover, our results also show the effectiveness of parameter-efficient fine-tuning: RepairLLaMA’s repair adapters, with only 4M parameters, are 1600x smaller than the initial pre-trained LLM (CodeLLama-7B). To sum up, the efficient representations and repair adapters of RepairLLaMA outperform recent results on fine-tuning for program repair (Jiang et al., 2023; Huang et al., 2023a; Wang et al., 2023b) as well as world-class models such as GPT-3.5 and GPT-4.
To sum up, we make the following contributions:
@@
  • We design RepairLLaMA, an original fine-tuning pipeline for automated program repair, that maximizes knowledge from the program repair domain, while keeping strong alignment with pre-training.
@@
  • We systematically evaluate code representations for program repair fine-tuning. Our results clearly show that the best code representation leverages task-specific signals, including fault localization and original buggy code.
@@
  • We demonstrate that parameter-efficient fine-tuning performs better than full-parameter fine-tuning in the context of program repair. The “repair adapters” of RepairLLaMA are training-efficient, inference-efficient, and powerful to repair bugs, achieving state-of-the-art performance across two benchmarks, Defects4J and HumanEval-Java, outperforming even GPT-4.
@@
  • For the sake of open science, we publish our source code, models, and artifacts at https://github.com/ASSERT-KTH/repairllama and provide a demo website at https://repairllama.github.io.
@@
Figure 1. Overview of RepairLLaMA. The core novelties of RepairLLaMA are the APR specific code representations and the engineering of an effective program repair adapter that is plugged into the underlying LLM.
2. RepairLLaMA: Efficient Fine-Tuning for Program Repair
2.1. Overview
@@
Figure 1 illustrates the pipeline of RepairLLaMA for APR, which is divided into three consecutive stages. The core novelties of this pipeline are: 1) the APR specific code representations and 2) the end-to-end use of a parameter-efficient fine-tuning technique.
@@
2.2. Target Bugs
The first consideration when designing a fine-tuning pipeline for program repair is the bugs we aim to fix. This relates to 1) the programming language, 2) the type of bugs (syntax errors, runtime errors, functional errors, etc), and, 3) the difficulty of bugs, which can be proxied by the code span to modify in order to fix the bug.
In this work, we focus on 1) Java bugs, 2) that are functional, and come with at least a failing test case, and, 3) that are intra-procedural with any length of the code span, i.e. that can be fixed with changes to a single function (called hereafter single-function bugs).
Note that we do want to support bugs that require changes in multiple locations in the function (Ye and Monperrus, 2023), beyond single-line or single-chunk bugs.
2.3. Choice of the Initial LLM
@@
First, the LLM should be publicly available and open-source. Fine-tuning a closed-source LLM on the task-specific dataset is not a valid option. Although some companies like OpenAI do provide an API for fine-tuning their LLMs, it is expensive, and the ownership of the final model (incl. weights) does not meet open-science reproduction criteria. Open-source models, such as LLaMA (Touvron et al., 2023) or StarCoder (Li et al., 2023), publish model weights online, allowing anyone to modify and deploy them.
Second, the LLM should be pre-trained with large-scale code data. As observed by related work (Li et al., 2023; Roziere et al., 2023), LLMs pre-trained on massive code data achieve better performance in code-related tasks. Thus, we consider only LLMs specialized on code.
Third, the initial LLM should have been trained with an infilling objective (Bavarian et al., 2022) during pre-training. As observed by related work (Jiang et al., 2023), infilling is a natural and effective learning objective for the program repair task, since it allows the model to synthesize code according to both the context appearing before and after. It should also be supported by an off-the-shelf parameter-efficient fine-tuning library.
@@
2.4. Choice of Code Representations
Source code representation is a critical aspect that significantly impacts the effectiveness of the model (Namavar et al., 2022). In this section, we discuss key characteristics of the source code representation design space. We introduce, motivate, and elaborate on input and output code representations specific to the program repair task.
2.4.1. Representation of Fault Localization
Virtually all the APR literature assumes line-based fault localization, with a single line given as input to the repair algorithm. This is not appropriate to fix multi-location bugs (Saha et al., 2019; Ye and Monperrus, 2023). Consider Figure 3 (OR4), which shows the canonical patch for the multi-location bug Chart-5 from Defects4J. In this case, fault localization must identify a location where an entirely new if block should be synthesized and inserted as well as another pre-existing if condition, appearing later in the code. To our knowledge, there is no fault localization technique able to predict tuples of blocks to be repaired together.
@@ 2.4.1. Representation of Fault Localization
2.4.2. Input Representation Space
Figure 2. Buggy code of the multi-location bug Chart-5 represented in our four different input representations.
@@ 2.4.2. Input Representation Space
IR3: Buggy function w/ infilling mask This representation uses the infilling scheme some LLMs are trained for during pre-training (Bavarian et al., 2022). The buggy chunk is replaced by the infilling token, which prompts the model to fill it. For example, in Figure 2 (IR3), the three lines between the start and end of the suspicious region are replaced by the ¡FILL_ME¿ token. This representation yields shorter inputs and requires less fine-tuning since the infilling objective has been used during pre-training. However, by masking the buggy portion of code, this representation incurs information loss that can be useful to generate a fix.
@@
2.4.3. Output Representation Space
Figure 3. Patch for multi-location bug Chart-5 represented in our four different output representations.
@@ 2.4.3. Output Representation Space
2.4.4. Input/Output Representation Pairs
Table 1. Possible code representation pairs for fine-tuning LLMs for automated program repair. They exploit the characteristics of the APR task, incl. the presence of fault localization signals and the notion of “buggy code”.
@@ 2.4.4. Input/Output Representation Pairs
To utilize an LLM for APR, input and output representations must be carefully paired. This is because all input representations cannot be paired with all output representations. For instance, IR1 cannot pair with OR2 since one cannot apply a fixed chunk to the buggy function without the fault localization information. Table 1 provides the list of the code representation pairs that are studied in this paper. Each row corresponds to a code representation pair. Column FL indicates whether the pair includes or not fault localization information. Column Aligned w/ PT provides a relative assessment of the alignment of the representation w.r.t. the pre-training data/objective. A red cross means that the code representation is not aligned with the pre-training data and objective. The left side shows the input and the right the output representations. Column Buggy Code indicates whether the pair includes or not the original buggy code.
@@
The latter three rows (i.e., IR2xOR2, IR3xOR2, IR4xOR2) include code representation pairs that contain fault localization information, either as tokens or as infilling, which is specific to program repair. The most aligned representation with pre-training is IR3xOR2 since the pre-trained model has support for infilling. IR2 represents the infilling objective with never-before-seen comments, whereas IR4 keeps the buggy code as comments. The natural output representation to pair with these is OR2 since it only includes the new code to replace the already localized buggy chunk, minimizing output length. Note that we have empirically tested other combinations in a pilot experiment, and the ones not listed in Table 1 underperform.
2.5. Choice of Fine-Tuning Dataset
@@
2.6. Fine-Tuning Repair Adapters with LoRA
With the recent release of various LLMs, the scale of parameters has significantly increased. For instance, state-of-the-art models such as LLaMA (Lu et al., 2023) and CodeLLaMA (Roziere et al., 2023) range from 7B to 70B parameters. Fine-tuning these LLMs often requires substantial GPU resources. As an example, Lv et al. (Lv et al., 2023) report that fine-tuning the full parameters of LLaMA-7B on an RTX 3090 consumes 126.08 GB at peak GPU memory usage, with the batch size and sequence length set to 1 and 1024 respectively. Fine-tuning current LLMs with limited resources is a challenge.
RepairLLaMA uses LoRA (Hu et al., 2021), a state-of-the-art parameter-efficient fine-tuning method. LoRA freezes the LLM and fine-tunes low-rank matrices in each layer of the language model that are trainable. The trained matrices compose an “adapter”, which is many orders of magnitude smaller than the language model itself. In RepairLLaMA, the repair adapter is a LoRA adapter, dedicated to the program repair task.
2.7. Inference
The final step is to deploy the repair adapter. The target buggy program is fed to a fault localization algorithm and processed to generate an APR-specific code representation. Then, the code representation is fed to the initial model combined with the LoRA repair adapter, to generate a list of candidate patches for the buggy program at hand. Patches are then checked for plausibility and correctness per off-the-shelf techniques.
3. Experimental Methodology
3.1. Research Questions

### v5

@@
Abstract
Automated Program Repair (APR) has evolved significantly with the advent of Large Language Models (LLMs). Fine-tuning LLMs for program repair is a recent avenue of research, with many dimensions which have not been explored. Existing work mostly fine-tune LLMs with naive code representations and does not scale to frontier models. To address this problem, we propose RepairLLaMA, a novel program repair approach that 1) identifies optimal code representations for APR with fine-tuned models, and 2) pioneers state-of-the-art parameter-efficient fine-tuning technique (PEFT) for program repair. This results in RepairLLaMA producing a highly effective ‘program repair adapter’ for fixing bugs with AI. Our experiments demonstrate the validity of both concepts. First, fine-tuning adapters with program repair specific code representations enables the model to use meaningful repair signals and produce better patches. Second, parameter-efficient fine-tuning helps fine-tuning to converge and clearly contributes to the effectiveness of RepairLLaMA in fixing bugs outside the fine-tuning data distribution. Overall, RepairLLaMA correctly fixes 144 Defects4J v2, 109 HumanEval-Java, and 20 GitBug-Java bugs, outperforming all baselines.
Index Terms:
@@ Abstract
1 Introduction
@@
Automated program repair (APR) [1, 2] aims at automatically fixing a software bug without human intervention. Learning-based repair [3, 4, 5, 6, 7, 8, 9, 10] has become the mainstream solution to this problem due to the powerful ability of deep neural networks to learn complex bug fix patterns. Large language models (LLMs), pre-trained on vast amounts of data, have pushed learning-based repair to the next frontier [8, 11]. There are two main research lines in applying LLMs to program repair: (1) fine-tuning techniques to adapt pre-trained LLMs and create specialized repair models [5, 12, 8, 10], and (2) prompt engineering and agent-based approaches that leverage LLMs’ capabilities through carefully designed inputs and interaction patterns [13, 14, 15, 16, 17].
Fine-tuning has the potential to learn domain-specific representations and leverage signals in order to generate high-quality patches. The drawback is that it requires substantial computational resources and high-quality training data. Prompt engineering and agent-based approaches mostly consist of ad hoc human-developed prompts and workflows to instruct models for specific tasks and are bounded by the limited size of context windows. Both lines make different trade-offs in applying LLMs to automated bug fixing. In this paper, we make an important contribution in the space of fine-tuning LLMs for program repair.
@@
Fine-tuning LLMs for program repair is complex. Early work simply refines the network weights based on additional fine-tuning data. However, this kind of fine-tuning is rather primitive and suffers from two significant drawbacks. First, fine-tuning is also known to be able to adapt the input/output representations of the data under study [18]. In the context of program repair, there is an opportunity to fine-tune with code representations that maximize downstream task performance, that is, repair performance. In particular, previous work overlooks the realistic representation of fault localization in the input. Second, previous work considered the most basic fine-tuning technique, which is full-parameter fine-tuning. As LLMs increase in size [19], full-parameter fine-tuning poses important overfitting problems when fine-tuning data is limited, which is typically the case in program repair. In this paper, we address the problem of devising efficient fine-tuning techniques [20] for program repair, with a focus on code representations and adapters.
@@
We propose RepairLLaMA, a new program repair approach that leverages parameter-efficient fine-tuning to adapt LLMs to handle repair-specific code representations. First, RepairLLaMA’s code representations incorporate fault localization signals and are designed to support multi-location bugs. Second, RepairLLaMA utilizes Low-Rank Adaption (LoRA), a state-of-the-art parameter-efficient fine-tuning technique, to train a much smaller repair adapter (when compared to the full LLM) that adapts the LLM for program repair while helping prevent overfitting [21]. As we will demonstrate in this paper, the concept of repair adapter is novel and potent.
@@
Our experimental results validate RepairLLaMA’s core design. First, RepairLLaMA achieves state-of-the-art fine-tuning performance in three benchmarks, correctly fixing 144 Defects4J v2 [22] bugs, and 109 and 20 bugs, respectively, on recently proposed benchmarks HumanEval-Java [8] and GitBug-Java [23], which boosts internal and external validity. The experiments show that the devised code representations with repair signals allow the LLM to synthesize patches more effectively than the naive code-only representations. Also, RepairLLaMA clearly outperforms non-fine-tuned baselines, incl. GPT-4. Moreover, our results also show the effectiveness of parameter-efficient fine-tuning: RepairLLaMA’s repair adapters, with only 4M parameters, are 1600x smaller than the initial pre-trained LLM, CodeLLama-7B [24]. To sum up, the efficient representations and repair adapters of RepairLLaMA outperform recent results on fine-tuning for program repair [8, 25, 10] as well as world-class models such as GPT-3.5 and GPT-4.
Overall, we make the following contributions:
@@
  • We design RepairLLaMA, an original fine-tuning pipeline for automated program repair with LLMs. RepairLLaMA’s representations maximize knowledge from the program repair domain, while keeping strong alignment with pre-training.
@@
  • We systematically evaluate different code representations for program repair fine-tuning. Our results clearly show that the best code representation leverages the task-specific signals of fault localization and original buggy code.
@@
  • We demonstrate that parameter-efficient fine-tuning performs better than full-parameter fine-tuning in the context of program repair. The “repair adapters” of RepairLLaMA are training-efficient, and achieve state-of-the-art repair performance across three benchmarks, Defects4J, HumanEval-Java, and GitBug-Java, outperforming even GPT-4.
@@
  • For the sake of open science, we publish our source code, models, and artifacts at https://github.com/ASSERT-KTH/repairllama.
@@
Figure 1: Overview of RepairLLaMA. The core novelties of RepairLLaMA are the APR specific code representations and the engineering of an effective program repair adapter that is plugged into the underlying LLM.
2 RepairLLaMA: Efficient Fine-Tuning for Program Repair
2.1 Overview
@@
Figure 1 illustrates the pipeline of RepairLLaMA for APR, which is divided into three consecutive stages. The core novelties of this pipeline are: 1) the APR specific code representations, and 2) the end-to-end use of a parameter-efficient fine-tuning technique.
@@
2.2 Target Bugs
The first consideration when designing a fine-tuning pipeline for program repair is the bugs we aim to fix. This relates to 1) the programming language, 2) the type of bugs (syntax errors, runtime errors, functional errors, etc), and 3) the difficulty of bugs, which can be proxied by the code span to modify in order to fix the bug.
In this work, we focus on 1) Java bugs, 2) that are functional, and come with at least a failing test case, 3) that are intra-procedural, i.e. fixed with changes to a single function (called hereafter single-function bugs). We do not make any assumption on the length of the method under repair, and 4) explicitly support bugs that require changes in multiple locations in the function [26], beyond single-line or single-chunk bugs.
2.3 Choice of the Initial LLM
@@
First, the LLM should be publicly available and open-source. Fine-tuning a closed-source LLM on the task-specific dataset is not a valid option. Although some companies like OpenAI do provide an API for fine-tuning their LLMs, it is expensive, and the ownership of the final model (incl. weights) does not meet open-science reproduction criteria. Open-source models, such as LLaMA [27] or StarCoder [28], publish model weights online, allowing anyone to modify and deploy them.
Second, the LLM should be pre-trained with large-scale code data. As observed by related work [28, 24], LLMs pre-trained on massive code data achieve better performance in code-related tasks. Thus, we consider only LLMs specialized on code.
Third, the initial LLM should have been trained with an infilling objective [29] during pre-training. As observed by related work [8], infilling is a natural and effective learning objective for the program repair task, since it allows the model to synthesize code according to both the context appearing before and after. It should also be supported by an off-the-shelf parameter-efficient fine-tuning library.
@@
2.4 Choice of Code Representations
Source code representation is a critical aspect that significantly impacts the effectiveness of the model [30]. In this section, we discuss key characteristics of the source code representation design space. We introduce, motivate, and elaborate on input and output code representations specific to the program repair task.
2.4.1 Representation of Fault Localization
Virtually all the APR literature assumes line-based fault localization, with a single line given as input to the repair algorithm. This is not appropriate to fix multi-location bugs [31, 26]. Consider Figure 3 (OR4), which shows the canonical patch for the multi-location bug Chart-5 from Defects4J. In this case, fault localization must identify a location where an entirely new if block should be synthesized and inserted as well as another pre-existing if condition, appearing later in the code. To our knowledge, there is no fault localization technique able to predict tuples of blocks to be repaired together.
@@ 2.4.1 Representation of Fault Localization
2.4.2 Input Representation Space
Figure 2: Buggy code of the multi-location bug Chart-5 represented in our four different input representations.
@@ 2.4.2 Input Representation Space
IR3: Buggy function w/ infilling mask This representation uses the infilling scheme some LLMs are trained for during pre-training [29]. The buggy chunk is replaced by the infilling token, which prompts the model to fill it. For example, in Figure 2 (IR3), the three lines between the start and end of the suspicious region are replaced by the <FILL_ME> token. This representation yields shorter inputs and requires less fine-tuning since the infilling objective has been used during pre-training. However, by masking the buggy portion of code, this representation incurs information loss that can be useful to generate a fix.
@@
2.4.3 Output Representation Space
Figure 3: Patch for multi-location bug Chart-5 represented in our four different output representations.
@@ 2.4.3 Output Representation Space
2.4.4 Input/Output Representation Pairs
Table I: Possible code representation pairs for fine-tuning LLMs for automated program repair. They exploit the characteristics of the APR task, incl. the presence of fault localization signals and the notion of “buggy code”.
@@ 2.4.4 Input/Output Representation Pairs
To utilize an LLM for APR, input and output representations must be carefully paired. This is because all input representations cannot be paired with all output representations. For instance, IR1 cannot pair with OR2 since one cannot apply a fixed chunk to the buggy function without the fault localization information. Table I provides the list of the code representation pairs that are studied in this paper. Each row corresponds to a code representation pair. Column FL indicates whether the pair includes or not fault localization information. Column Aligned w/ PT provides a relative assessment of the alignment of the representation w.r.t. the pre-training data/objective. A red cross means that the code representation is not aligned with the pre-training data and objective. The left side shows the input and the right the output representations. Column Buggy Code indicates whether the pair includes or not the original buggy code.
@@
The latter three rows (i.e., IR2xOR2, IR3xOR2, IR4xOR2) include code representation pairs that contain fault localization information, either as tokens or as infilling, which is specific to program repair. The most aligned representation with pre-training is IR3xOR2 since the pre-trained model has support for infilling. IR2 represents the infilling objective with never-before-seen comments, whereas IR4 keeps the buggy code as comments. The natural output representation to pair with these is OR2 since it only includes the new code to replace the already localized buggy chunk, minimizing output length. Note that we have empirically tested other combinations in a pilot experiment, and the ones not listed in Table I underperform.
2.5 Choice of Fine-Tuning Dataset
@@
2.6 Program Repair Adapters for LLMs
With the recent release of various LLMs, the scale of parameters has significantly increased. For instance, state-of-the-art models such as LLaMA [19] and CodeLLaMA [24] range from 7B to 70B parameters. Fine-tuning these LLMs often requires substantial GPU resources. As an example, Lv et al. [32] report that fine-tuning the full parameters of LLaMA-7B on an RTX 3090 consumes 126.08 GB at peak GPU memory usage, with the batch size and sequence length set to 1 and 1024 respectively. Fine-tuning current LLMs with limited resources is a challenge.
RepairLLaMA uses LoRA [20], a state-of-the-art parameter-efficient fine-tuning method that reduces memory requirements while maintaining model performance. Instead of fine-tuning all model parameters, LoRA freezes the pre-trained LLM weights and injects trainable low-rank matrices into specific attention layers. In addition to reducing memory requirements, by reducing the number of trainable parameters LoRA also acts as a regularizer [21] helping prevent overfitting during fine-tuning.
In our implementation, we apply LoRA to the query (q_proj) and value (v_proj) projection matrices in each transformer’s self-attention layer. These projection matrices transform the input embeddings into query and value vectors used in the attention. For each target weight matrix $A\in\mathbb{R}^{d\times k}$, LoRA decomposes the update into a product of two lower-rank matrices $B\in\mathbb{R}^{d\times r}$ and $C\in\mathbb{R}^{r\times k}$, where $r$ is the rank. During inference, the effective weight matrix is $W=W_{0}+BA$, where $W_{0}$ is the frozen pre-trained weight.
The trained matrices can be interpreted as a repair adapter that is much smaller than the original LLMs. For example, with rank r=8, our adapter only requires training 0.39% [^1] Take CodeLLaMA-7B as an example, its hidden size is 4096, so the original parameter of a matrix is 4096 $\times$ 4096. As for the LoRA adapter, this value is 4096 $\times$ 8 + 4096 $\times$ 8. of the parameters compared to full fine-tuning. This parameter-efficient approach allows RepairLLaMA to achieve strong program repair performance while being trainable on a single GPU.
2.7 Inference Time
The final step is to deploy the repair adapter. The target buggy program is fed to a fault localization algorithm and processed to generate an APR-specific code representation. Then, the code representation is fed to the initial model combined with the LoRA repair adapter to generate a list of candidate patches for the buggy program. Patches are then checked for plausibility and correctness per off-the-shelf techniques.
3 Experimental Methodology
3.1 Research Questions